Solve the system of
equations by graphing
x+y=8
x-y=6
where do I start?

Answer by lwsshak3(11628)

Solve the system of
equations by graphing
x+y=8
x-y=6
**
rewrite equations into standard form for straight lines: y=mx+b, m=slope, b=y-intercept
x+y=8
x-y=6
..
y=-x+8
y=x-6
equate two equations
-x+8=x-6
2x=14
x=7
y=x-6=7-6=1
point of intersection:(7,1)
..
Graph the two lines above using the slope and y-intercepts or plot points and check if they cross at the point of intersection. They should also be perpendicular to each other because the slope of one line(1) is the negative reciprocal(-1) of the other